OpenStudy (anonymous):

Given the expression (8 + 6i) − (−2 + 7i), perform the indicated operation and write the answer in the form a + bi. 6 + 13i 12 − i 10 + 13i 10 − i

OpenStudy (er.mohd.amir):

8+6i+2-7i=10-i

jagr2713 (jagr2713):

@Er.Mohd.AMIR its better to guide the person to the answer instead of doing the work for them.

OpenStudy (er.mohd.amir):

ok

OpenStudy (er.mohd.amir):

if -(a-b)=-a+b after open bracket sing change if - before bracket.

OpenStudy (xapproachesinfinity):

First open the parenthesis like @Er.Mohd.AMIR did \[(8+6i)-(-2+7i)=8+6i+2-7i \longleftarrow ~here~we ~are~distributing~~-1~to ~(-2+7i)\]

OpenStudy (xapproachesinfinity):

to (-2+7i) the part that is not appearing in my replay

OpenStudy (xapproachesinfinity):

then juct collect like term and add up

OpenStudy (xapproachesinfinity):

Clear!